# Comparing Quantum Annealing and Spiking Neuromorphic Computing for Sampling Binary Sparse Coding QUBO Problems

### v1

@@ 3.1 Binary Sparse Coding on 128 variables with QEMC
Both Figure 4 and 5 show that iterating reverse annealing, applied to these fully connected QUBOs, significantly improves the found solutions. In particular, when a good anneal fraction is used for the reverse anneal pause, the solution quality monotonically improves. This is is significant to see for such large minor embedded optimization problems. The effectiveness of iterated reverse annealing, or QEMC, has not been empirically demonstrated for minor embedded systems of this size, as far as we are aware.
Like Figure 4, Figure 5 plots the best solution found when executing simulated annealing on the problem QUBOs - and like Figure 4, the simulated annealing solutions are also the optimal solutions found by CPLEX. In other words, the dashed horizontal black line in Figure 5 denotes the optimal energy of, or cost value, of these $128$ variable QUBOs.
The best anneal fraction, which tradeoffs between exploration of new solutions and retaining memory of the initial spin state, is usually either $0.5$ or $0.55$ and this is true for both Figure 4 and 5. Notably, using a reverse annealing pause at $s=0.4$ causes the Monte Carlo chain to not converge, but rather to oscillate approximately at the energy of the starting state at the beginning of the process.
@@ 4 Discussion
In this article we have demonstrated that moderately sized binary sparse coding QUBO models can be sampled using both quantum annealing and neuromorphic computing technologies. These binary sparse coding QUBO problems were constructed using an un-supervised and un-normalized dictionary learning approach. We have found that the iterated reverse quantum annealing technique, implemented on a D-Wave quantum annealer, performed better than standard forward annealing. Similarly, iterated warm starting on Loihi 2 improved the solution quality of sampling the QUBOs. However, this energy improvement was marginal. We also found that the Loihi 2 sampling was better than the sampling done on Loihi 1 [henke2023sampling]. Notably, a consistent trend for the Loihi processor sampling is that the generated samples tend to be quite sparse – more sparse than sub-optimal sampling from standard simulated annealing or quantum annealing.
Amongst others, the following three reasons might have impeded the forward annealing and reverse annealing sampling from reaching the global minimum in our experiments.
  1. 1.
    The first, and almost certainly the most important, is the coefficient precision limits that can be programmed on current D-Wave devices. The sparse coding QUBOs have quadratic terms which are defined with high precision, and are nearly all very close to zero (see Figure 1). It seems very likely that such sparse coding QUBOs are simply not able to be correctly programmed at the hardware level due to the current limitations of the device.
  2. 2.
    The next obstacle is the use of a minor embedding. A minor embedding is necessary for implementing general QUBO problems on D-Wave, however such embeddings can have ferromagnetic chains which dominate the energy scale programmed on the chip, therefore not allowing a larger range of the logical problem coefficients to be effectively used in the computation. Additionally, although majority vote unembedding is used, chain breaks that are present in the samples represent likely sub-optimal solution quality.
  3. 3.
    Interaction with the environment that can cause biased noise.
@@
For future research, it would be interesting to be able to formulate sparse coding QUBO problems that match a desired connectivity graph. This would be particularly useful for generating hardware-compatible QUBOs for D-Wave hardware graphs, as it would allow one to considerably scale up the size of QUBO problems that can be solved (up to the size of the entire hardware graph, which has over $5000$ qubits). In particular, probing the scaling of solution quality for future generations of both neuromorphic processors and quantum annealing processors would be of interest to compare against existing classical methods.
@@ Appendix A Parallel Minor Embeddings
Figure 6: Seven disjoint minor embeddings of size $64$ variables with all-to-all connectivity (left), and two disjoint minor embeddings of size $128$ variables with all-to-all connectivity (right) on the Pegasus $P_{16}$ hardware graph of the Advantage_system4.1.
Figure 7: Schedule calibration for the exact $A(s)$ and $B(s)$ values in terms of GHz as a function of the anneal fraction $s\in[0,1]$.
@@ Appendix B D-Wave Quantum Annealer Schedule Calibration
Figure 7 gives the exact vendor-provided anneal schedule calibration data.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy through Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (under the contract No. 89233218CNA000001). The research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under the project number 20220656ER. This research used resources (D-Wave Leap Cloud access) provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Research presented in this article was supported by the NNSA’s Advanced Simulation and Computing Beyond Moore’s Law Program at Los Alamos National Laboratory. We gratefully acknowledge support from the Advanced Scientific Computing Research (ASCR) program office in the Department of Energy’s (DOE) Office of Science, award #77902. This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This paper has been assigned LANL report number LA-UR-23-26361

### v2

@@ 3.1 Binary Sparse Coding on 128 variables with QEMC
Both Figure 5 and 6 show that iterating reverse annealing, applied to these fully connected QUBOs, significantly improves the found solutions. In particular, when a good anneal fraction is used for the reverse anneal pause, the solution quality monotonically improves. This is is significant to see for such large minor embedded optimization problems. The effectiveness of iterated reverse annealing, or QEMC, has not been empirically demonstrated for minor embedded systems of this size, as far as we are aware.
Like Figure 5, Figure 6 plots the best solution found when executing simulated annealing on the problem QUBOs; and like Figure 5, the simulated annealing solutions are also the optimal solutions found by CPLEX. In other words, the dashed horizontal black line in Figure 6 denotes the optimal energy of, or cost value, of these $128$ variable QUBOs.
The best anneal fraction, which tradeoffs between exploration of new solutions and retaining memory of the initial spin state, is usually either $0.5$ or $0.55$ – this is true for both Figure 5 and Figure 6. Notably, using a reverse annealing pause at $s=0.4$ causes the Monte Carlo chain to not converge, but rather to oscillate approximately at the energy of the starting state at the beginning of the process.
@@ 4 Discussion
In this article we have demonstrated that moderately sized binary sparse coding QUBO models can be sampled using both quantum annealing and neuromorphic computing technologies. These binary sparse coding QUBO problems were constructed using an un-supervised and un-normalized dictionary learning approach. We have found that the iterated reverse quantum annealing technique, implemented on a D-Wave quantum annealer, performed better than standard forward annealing. Similarly, iterated warm starting on Loihi 2 improved the solution quality of sampling the QUBOs. However, this energy improvement was marginal. We also found that the Loihi 2 sampling was better than the sampling done on Loihi 1 [henke2023sampling]. Notably, a consistent trend for the Loihi processor(s) sampling is that the generated samples tend to be quite sparse – more sparse than sub-optimal sampling from standard simulated annealing or quantum annealing.
Amongst others, the following reasons might have impeded the forward annealing and reverse annealing sampling from reaching the global minimum in our experiments. The first, and almost certainly the most important, is the coefficient precision limits that can be programmed on current D-Wave devices. The sparse coding QUBOs have quadratic terms which are defined with high precision, and are nearly all very close to zero (see Figure 1). It is very likely that such sparse coding QUBO coefficients require higher precision than what the D-Wave hardware supports. Note that the precision limits on the hardware are not a well defined bit precision, but rather are influenced by a number of factors on the analog hardware, such as asymmetric digital-to-analog converter quantization, making the exact precision limits hard to know for all problem types and simulation settings  [^5] https://docs.dwavesys.com/docs/latest/c_qpu_ice.html. The next obstacle is the use of a minor embedding. A minor embedding is necessary for implementing general QUBO problems on D-Wave, however such embeddings have ferromagnetic chains which dominate the energy scale programmed on the chip, therefore not allowing a larger range of the logical problem coefficients to be effectively used in the computation. Additionally, although majority vote unembedding is used, chain breaks that are present in the samples represent likely sub-optimal solution quality. Analog control errors when the hardware is representing the diagonal Hamiltonian on the physical machine can also lead to errors in the simulation. Finally, interaction with the environment causes the system to decohere and noise to be introduced into the simulation.
@@
For future research, it would be interesting to be able to formulate sparse coding QUBO problems that match a desired connectivity graph. This would be particularly useful for generating hardware-compatible QUBOs for D-Wave hardware graphs, as it would allow one to considerably scale up the size of QUBO problems that can be solved (up to the size of the entire hardware graph, which has over $5000$ qubits). In particular, probing the scaling of solution quality for future generations of both neuromorphic processors and quantum annealing processors would be of interest to compare against existing classical optimization algorithms. In general, creating significantly larger sparse coding QUBOs using un-normalized dictionary learning [henke2023sampling], and then evaluating methods to find good solutions to those larger QUBOs, would be of interest for future studies.
@@ Appendix A Parallel Minor Embeddings
Figure 7: Parallel quantum annealing fully connected minor-emebddings on the hardware graph of Advantage_system4.1. Seven disjoint minor embeddings of size $64$ variables with all-to-all connectivity (left), and two disjoint minor embeddings of size $128$ variables with all-to-all connectivity (right) on the Pegasus $P_{16}$ hardware graph of the Advantage_system4.1.
Figure 8: Calibrated anneal schedule for the D-Wave quantum annealing processor used in this study. The exact $A(s)$ and $B(s)$ values are plotted in units of GHz as a function of the anneal fraction $s\in[0,1]$.
@@ Appendix B D-Wave Quantum Annealer Schedule Calibration
Figure 8 gives the exact vendor-provided anneal schedule calibration data. This data is publicly available on the D-Wave website under QPU-Specific Characteristics  [^6] https://docs.dwavesys.com/docs/latest/doc_physical_properties.html.
@@ Acknowledgments
This work was supported by the U.S. Department of Energy through Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of the U.S. Department of Energy (under the contract No. 89233218CNA000001). The research presented in this article was supported by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under the project number 20220656ER and 20240032DR. This research used resources (D-Wave Leap Cloud access) provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. Research presented in this article was supported by the NNSA’s Advanced Simulation and Computing Beyond Moore’s Law Program at Los Alamos National Laboratory. We gratefully acknowledge support from the Advanced Scientific Computing Research (ASCR) program office in the Department of Energy’s (DOE) Office of Science, award #77902. This research used resources provided by the Darwin testbed at Los Alamos National Laboratory (LANL) which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This paper has been assigned LANL report number LA-UR-23-26361
Data Availability Statement
Data is publicly available on Zenodo [pelofske_2025_14735999].